OpenStudy (anonymous):

Find the value of x so that the line passing through the points will have the given slope: (x, 6) and (10, -3) with a slope of -3/2. A. -13 B. 4 C. 13 D. -12 Can someone explain how to do this? I came out with an answer of x=20, so obviously I'm doing something wrong.

OpenStudy (anonymous):

Use \[ m=\frac{y_2-y_1}{x_2-x_1} \]\[ -\frac 32 = \frac{-3-6}{10-x} \]

OpenStudy (anonymous):

I set up the equation correctly, I must have made a mistake within solving it. @wio

OpenStudy (anonymous):

Cross multiply next?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

I am confused on what to do with the \[- \frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

Cross multiply like you said

OpenStudy (anonymous):